# A $2$-distance $(2\Delta+7)$-coloring of planar graphs

### v2

@@ 2.2. The case $\Delta=7$
Recall that $G$ does not admit any $2$-distance $21$-coloring, but any planar graph $G^{\prime}$ obtained from $G$ with smaller $|V(G^{\prime})|+|E(G^{\prime})|$ admits a 2-distance $21$-coloring. We begin with some structural properties of $G$ similar to the case when $\Delta=6$.
The proof of Lemma 2.16 is omitted since it is similar to Lemma 2.1.
@@
If $v$ is a $3$-vertex, then $m_{3}(v)=0$, $m_{4}(v)\leq 1$ and each neighbour of $v$ is a $7(5^{-})$-vertex.
@@
Let $v$ be a $3$-vertex. Assume for a contradiction that $v$ is incident to a $3$-face, say $f_{1}=v_{1}vv_{2}$. It is clear that $d_{2}(v)\leq 19$. If we set $G^{\prime}=G-v+\{v_{1}v_{3}\}$, then $G^{\prime}$ would be proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $21$ colors. Since $d_{2}(v)\leq 19$, we can color $v$ with an available color, a contradiction.
Suppose now that $m_{4}(v)\geq 2$, i.e., $v$ is incident to two $4$-faces, say $f_{1}=v_{1}vv_{2}x$ and $f_{2}=v_{2}vv_{3}y$. Clearly, we have $d_{2}(v)\leq 19$. If we set $G^{\prime}=G-v+\{v_{1}v_{3}\}$, then $G^{\prime}$ would be proper with respect to $G$. Similarly as above, we get a contradiction.
Next we claim that each neighbour of $v$ is a $7(5^{-})$-vertex. Note that $v$ cannot have any $7(6^{+})$-neighbour since $m_{3}(v)=0$. Moreover, $v$ cannot have any $6^{-}$-neighbour. Indeed, if $v_{1}$ is such a vertex, then we set $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3}\}$. In such a case, $d_{2}(v)\leq 20$, and $G^{\prime}$ is proper with respect to $G$. Similarly as above, we get a contradiction. Thus, each neighbour of $v$ is a $7(5^{-})$-vertex. ∎
@@
If $v$ is a $4$-vertex, then $m_{3}(v)\leq 3$, i.e., $G$ has no 4(4)444(4)4 ( 4 )-vertex.
@@
Let $v$ be a $4$-vertex. Assume that $m_{3}(v)=4$, and so we have $d_{2}(v)\leq 20$. If we set $G^{\prime}=G-v$, then $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $21$ colors. Since $d_{2}(v)\leq 20$, we can color $v$ with an available color, a contradiction. ∎
Let us define some vertices having special forms: If $v$ is a $4(1)$-, $4(2)$- or $4(3)$-vertices, then we call $v$ as a bad $4$-vertex. Similarly, if $v$ is a $5(4)$- or 5(5)555(5)5 ( 5 )-vertices, then we call $v$ as a bad $5$-vertex.
@@
    If $m_{3}(v)\geq 1$, and $v$ has a $4$-neighbour, then $v$ cannot have any $5(4)$-neighbour.
@@
$(a)$ By contradiction, assume that $v$ has a 5(5)555(5)5 ( 5 )-neighbour. In such a case, $v$ should be either $4(2)$-vertex or $4(3)$-vertex, so we may assume that $vv_{2}$ is contained in two $3$-faces. Observe that $d_{2}(v)\leq 20$. If we set $G^{\prime}=G-v+\{v_{2}v_{4}\}$, then $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $21$ colors. Since $d_{2}(v)\leq 20$, we can color $v$ with an available color, a contradiction.
$(b)$ Let $m_{3}(v)\geq 1$. Suppose that $v$ has two $4$-neighbours, and let $v_{1}$ be one of those $4$-vertices. In this case, we have $d_{2}(v)\leq 20$. If we set $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3},v_{1}v_{4}\}$, then $G^{\prime}$ is proper with respect to $G$. In a similar way as above, we get a contradiction.
$(c)$ Let $m_{3}(v)\geq 1$. Assume that $v$ has both a $4$-neighbour and a $5(4)$-neighbour. In this case, we have $d_{2}(v)\leq 20$. Similarly as above, we get a contradiction. ∎
@@
Let $v$ be a $4(1)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{4}$ the faces incident to $v$ such that $f_{1}=v_{1}vv_{2}$.
$(a)$ Let $m_{4}(v)=2$. We first show that $v$ has a $7$-neighbour. Assume for a contradiction that all neighbours of $v$ are $6^{-}$-vertices. In such a case, we have $d_{2}(v)\leq 20$ since $m_{4}(v)=2$. If we set $G^{\prime}=G-v+\{v_{2}v_{3},v_{3}v_{4},v_{1}v_{4}\}$, then $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $21$ colors. Since $d_{2}(v)\leq 20$, we can color $v$ with an available color, a contradiction.
Now we assume that $v$ has no $7(5^{-})$-neighbour. Since $v$ has no any 7(7)777(7)7 ( 7 )-neighbour as $m_{3}(v)=1$, the vertex $v$ is incident to a $7(6)$-vertex, say $v_{i}$. Clearly we have $i=1$ or $i=2$, say $i=1$. If the other neighbours of $v$ are $6^{-}$-vertices, then we have $d_{2}(v)\leq 20$. When we set $G^{\prime}=G-v+\{v_{2}v_{3},v_{3}v_{4},v_{1}v_{4}\}$, the graph $G^{\prime}$ is proper with respect to $G$. Similarly as above, we get a contradiction. This infers that $v$ has one more $7(6)$-neighbour.
$(b)$ Let $m_{4}(v)=3$. We first show that $v$ has two $7$-neighbours. Assume for a contradiction that three neighbours of $v$ are $6^{-}$-vertices. In such a case, we have $d_{2}(v)\leq 20$ since $m_{4}(v)=3$. If we set $G^{\prime}=G-v+\{v_{2}v_{3},v_{1}v_{4}\}$, then $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $21$ colors. Since $d_{2}(v)\leq 20$, we can color $v$ with an available color, a contradiction. It then follows that $v$ has two $7(6^{-})$-neighbours, since $v$ cannot have any 7(7)777(7)7 ( 7 )-neighbour. ∎
@@
    If $m_{4}(v)=0$, then $v$ has either a $7(5^{-})$-neighbour and a $6(4^{-})$-neighbour or two $7(6^{-})$-neighbours.
@@
    If $m_{4}(v)=1$, then $v$ has either two $7(6^{-})$-neighbours and two $6(4^{-})$-neighbours or three $7(6^{-})$-neighbours.
@@
    If $m_{4}(v)=2$, then $v$ has either two $7(5^{-})$-neighbours and two $7(6)$-neighbours or three $7(5^{-})$-neighbours.
@@
Let $v$ be a $4(2)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{4}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ and $f_{j}=v_{j}vv_{j+1}$ with $i<j$ are $3$-faces (see Figure 3).
We remark that if $d_{2}(v)\leq 20$, then we set $G^{\prime}=G-v+\{v_{2}v_{4}\}$ (when $i=1$, $j=2$) or $G^{\prime}=G-v+\{v_{2}v_{3},v_{1}v_{4}\}$ (when $i=1$, $j=3$), and so the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $21$ colors. Since $d_{2}(v)\leq 20$, we can color $v$ with an available color, a contradiction. We therefore assume that $d_{2}(v)\geq 21$.
$(a)$ Let $m_{4}(v)=0$. Suppose first that $v$ has a 7(7)777(7)7 ( 7 )-neighbour. If $v$ has two $6^{-}$-neighbours, then $d_{2}(v)\leq 20$, a contradiction. This infers that $v$ has two $7(6^{-})$-neighbours. Now we assume that $v$ has no 7(7)777(7)7 ( 7 )-neighbour. Clearly, $v$ has at least one $7$-neighbour, since otherwise, we have $d_{2}(v)\leq 20$ when all neighbours of $v$ are $6^{-}$-vertices, a contradiction. So, $v$ has a $7(6^{-})$-neighbour, say $v_{r}$. Assume that $v$ has exactly one $7(6^{-})$-neighbour, since otherwise, the claim holds. So, the other neighbours of $v$ are $6^{-}$-vertices. If $v_{r}$ is a $7(6)$-vertex or $v$ has a $6(5^{+})$-neighbour, then we have again $d_{2}(v)\leq 20$, a contradiction. Thus, we conclude that $v_{r}$ is a $7(5^{-})$-vertex, as well as $v$ has a $6(4^{-})$-neighbour.
$v$$v_{2}$$v_{3}$$v_{4}$$v_{1}$
@@
Figure 3. Two possible $3$-faces $f_{i}=v_{i}vv_{i+1}$ and $f_{j}=v_{j}vv_{j+1}$ incident to a $4(2)$-vertex.
$(b)$ Let $m_{4}(v)=1$. We first claim that $v$ has at least two $7$-neighbours. Indeed, if three neighbours of $v$ are $6^{-}$-vertices, then we have $d_{2}(v)\leq 20$ as $m_{4}(v)=2$, a contradiction. Notice that $v$ has at most one 7(7)777(7)7 ( 7 )-neighbour as $m_{3}(v)=2$. We distinguish the remainder of the proof based on the number of edges $v_{i}v_{i+1}$, $v_{j}v_{j+1}$ contained in two $3$-faces.
Suppose that both $v_{i}v_{i+1}$ and $v_{j}v_{j+1}$ are contained in two $3$-faces. In this case, $v$ cannot have a $6^{-}$-neighbour, since otherwise, we have $d_{2}(v)\leq 20$, a contradiction. This infers that $v$ has three $7(6^{-})$-neighbours.
Next we suppose that exactly one of the edges $v_{i}v_{i+1}$, $v_{j}v_{j+1}$ is contained in two $3$-faces. In this case, $v$ cannot have any 7(7)777(7)7 ( 7 )- and 6(6)666(6)6 ( 6 )-neighbours, i.e., $v$ has two $7(6^{-})$-neighbours, say $v_{1},v_{2}$. Consider the vertices $v_{3},v_{4}$, if one of them is a $5^{-}$-vertex or both are $6$-vertices, then we have again $d_{2}(v)\leq 20$, a contradiction. Thus, one of $v_{3},v_{4}$ is a $7(6^{-})$-vertex while the other is a $6^{+}$-vertex. This implies that $v$ has one more $7(6^{-})$-neighbour, i.e., $v$ has totally at least three $7(6^{-})$-neighbours.
Suppose now that none of the edges $v_{i}v_{i+1}$, $v_{j}v_{j+1}$ is contained in two $3$-faces. In this way, $v$ cannot have any 7(7)777(7)7 ( 7 )-, $7(6)$-, 6(6)666(6)6 ( 6 )- and $6(5)$-neighbours. So, $v$ has two $7(5^{-})$-neighbours, say $v_{1},v_{2}$. Consider the vertices $v_{3},v_{4}$, if one of them is a $5^{-}$-vertex, then the other must be $7(5^{-})$-vertex, since otherwise we have $d_{2}(v)\leq 20$, a contradiction. If none of $v_{3},v_{4}$ is a $5^{-}$-vertex, then we deduce that either one of $v_{3},v_{4}$ is a $7(5^{-})$-vertex or both are $6(4^{-})$-vertices.
$(c)$ Let $m_{4}(v)=2$. Notice first that if $v$ has two $6^{-}$-neighbours or one 7(7)777(7)7 ( 7 )-neighbour or one 6(6)666(6)6 ( 6 )-neighbour, then we have $d_{2}(v)\leq 20$ since $m_{4}(v)=2$, a contradiction. Therefore, we deduce that $v$ has three $7(6^{-})$-neighbours. In fact, if $v$ has three $7(6)$-neighbours, then both $v_{i}v_{i+1}$ and $v_{j}v_{j+1}$ must be contained in two $3$-faces, and so we have again $d_{2}(v)\leq 20$, a contradiction. Hence, $v$ has two $7(6^{-})$-neighbours and one $7(5^{-})$-neighbours. Moreover, if $v$ has two $7(6)$-neighbours, the other neighbours of $v$ must be $7(5^{-})$-vertices, since otherwise, we have $d_{2}(v)\leq 20$, a contradiction. Consequently, the claim holds. ∎
@@
    If $m_{4}(v)=0$, then $v$ has either four $7(5^{-})$-neighbours or three $7(5^{-})$-neighbours and a $7(6)$-neighbour or three $7(5^{-})$-neighbours and a $6(4^{-})$-neighbour or two $7(5^{-})$-neighbours and two $7(6)$-neighbours.
@@
Let $v$ be a $4(3)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{4}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for each $i\in[3]$ is a $3$-face.
We first claim that at most one edge in $G[N(v)]$ is contained in two $3$-faces. Indeed, if there exist such two edges, then we would have $d_{2}(v)\leq 20$. When we set $G^{\prime}=G-v+\{v_{1}v_{4}\}$, the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $21$ colors. Since $d_{2}(v)\leq 20$, we can color $v$ with an available color, a contradiction. This particularly infers that $v$ has neither 6(6)666(6)6 ( 6 )- nor 7(7)777(7)7 ( 7 )-neighbour. In a similar way, we deduce that $v$ cannot have any $6(5)$-neighbour, two $6(4^{-})$-neighbours, and three $7(6)$-neighbours.
$(a)$ Let $m_{4}(v)=0$. As earlier stated, $v$ has at most one $6(4^{-})$-neighbours. In fact, if $v$ has a $6(4^{-})$-neighbour, then $v$ cannot have any $7(6)$-neighbours, since otherwise, we have $d_{2}(v)\leq 20$, and similarly as above, we get a contradiction. This implies that the neighbours of $v$ other than $6(4^{-})$-vertex are $7(5^{-})$-vertices. Suppose now that $v$ has no $6(4^{-})$-neighbour. Since $v$ has at most two $7(6)$-neighbours, we have either four $7(5^{-})$-neighbours or three $7(5^{-})$-neighbours and a $7(6)$-neighbour or two $7(5^{-})$-neighbours and two $7(6)$-neighbours.
$(b)$ Let $m_{4}(v)=1$. If $v$ has a $7(6^{+})$- or $6^{-}$-neighbour, then $d_{2}(v)\leq 20$, and similarly as above, we get a contradiction. ∎
@@
    $v$ cannot have two 5(5)555(5)5 ( 5 )-neighbours.
@@
    If $v$ has a 5(5)555(5)5 ( 5 )-neighbour, then $v$ cannot be adjacent to a $4$-vertex.
@@
    If $v$ has a $7(6)$-neighbour, then $v$ cannot be adjacent to two $4$-vertices.
@@
    If $m_{4}(v)=0$, then $v$ has two neighbours different from $5^{-}$- and 6(6)666(6)6 ( 6 )-vertex.
@@
    If $m_{4}(v)=1$, then $v$ has at least three $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex.
@@
    If $m_{4}(v)=1$, and $v$ has a 7(7)777(7)7 ( 7 )-neighbour, then $v$ has at least four $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex.

### v3

@@ 2.2. The case $\Delta=7$
Recall that $G$ does not admit any $2$-distance 2121-coloring, whereas any planar graph $G^{\prime}$ obtained from $G$ with a smaller value of $|V(G^{\prime})|+|E(G^{\prime})|$ admits a $2$-distance 2121-coloring. We begin by establishing several structural properties of $G$, similar to those in the case $\Delta=6$.
The proof of Lemma 2.18 is omitted, as it is similar to the proof of Lemma 2.3.
@@
If $v$ is a $3$-vertex, then $m_{3}(v)=0$, $m_{4}(v)\leq 1$, and each neighbour of $v$ is a $7(5^{-})$-vertex.
@@
Let $v$ be a $3$-vertex. Assume first that $v$ is incident to a $3$-face, say $f_{1}=v_{1}vv_{2}$. By Lemma 2.2, we have $d_{2}(v)\leq 19$. Setting $G^{\prime}=G-v+\{v_{1}v_{3}\}$ yields a graph that is proper with respect to $G$, contradicting Remark 2.1. Thus $m_{3}(v)=0$.
Next, suppose that $m_{4}(v)\geq 2$, i.e., $v$ is incident to two $4$-faces, say $f_{1}=v_{1}vv_{2}x$ and $f_{2}=v_{2}vv_{3}y$. Again, Lemma 2.2 gives $d_{2}(v)\leq 19$. The graph $G^{\prime}=G-v+\{v_{1}v_{3}\}$ is proper with respect to $G$, giving the same contradiction. Hence $m_{4}(v)\leq 1$.
Finally, we show that each neighbour of $v$ must be a $7(5^{-})$-vertex. Since $m_{3}(v)=0$, the vertex $v$ cannot have a $7(6^{+})$-neighbour. Moreover, $v$ cannot have any $6^{-}$-neighbour; otherwise, if $v_{1}$ is such a neighbour, then considering $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3}\}$ gives $d_{2}(v)\leq 20$ by Lemma 2.2, and $G^{\prime}$ is proper with respect to $G$, again a contradiction. Thus all neighbours of $v$ are $7(5^{-})$-vertices. ∎
@@
If $v$ is a $4$-vertex, then $m_{3}(v)\leq 3$. In particular, $G$ has no $4(4)$-vertex.
@@
Let $v$ be a $4$-vertex. Assume that $m_{3}(v)=4$. Then $d_{2}(v)\leq 20$ by Lemma 2.2. If we take $G^{\prime}=G-v$, then $G^{\prime}$ is proper with respect to $G$, contradicting Remark 2.1. Thus $m_{3}(v)\leq 3$. ∎
We now introduce some vertices of special types: a $4(1)$-, $4(2)$-, or $4(3)$-vertex is called a bad $4$-vertex, and a $5(4)$- or $5(5)$-vertex is called a bad $5$-vertex.
@@
    If $m_{3}(v)\geq 1$ and $v$ has a $4$-neighbour, then $v$ cannot have any $5(4)$-neighbour.
@@
$(a)$ Suppose, for a contradiction, that $v$ has a $5(5)$-neighbour, say $v_{2}$. We may therefore assume that each of $v_{1}v_{2},v_{2}v_{3}$ is contained in two $3$-faces. By Lemma 2.2, we have $d_{2}(v)\leq 20$. Setting $G^{\prime}=G-v+\{v_{2}v_{4}\}$ yields a graph that is proper with respect to $G$, contradicting Remark 2.1.
$(b)$ Let $m_{3}(v)\geq 1$. Suppose that $v$ has two $4$-neighbours, and let $v_{1}$ be one of them. Then $d_{2}(v)\leq 20$ by Lemma 2.2. Taking $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3},v_{1}v_{4}\}$ gives a graph that is proper with respect to $G$, a contradiction.
$(c)$ Let $m_{3}(v)\geq 1$. Suppose that $v$ has both a $4$-neighbour and a $5(4)$-neighbour. Then Lemma 2.2 again gives $d_{2}(v)\leq 20$, and the same reduction as in part (b) contradicts Remark 2.1. ∎
@@
Let $f_{1}=v_{1}vv_{2}$.
$(a)$ Suppose $m_{4}(v)=2$. We first show that $v$ has a $7$-neighbour. Assume all neighbours of $v$ are $6^{-}$-vertices. Then $d_{2}(v)\leq 20$ by Lemma 2.2. Setting $G^{\prime}=G-v+\{v_{2}v_{3},v_{3}v_{4},v_{1}v_{4}\}$ gives a graph that is proper with respect to $G$, contradicting Remark 2.1. Thus $v$ has at least one $7$-neighbour.
Now assume $v$ has no $7(5^{-})$-neighbour. Since $m_{3}(v)=1$, $v$ cannot have any $7(7)$-neighbour, so it must have a $7(6)$-neighbour. If the remaining neighbours of $v$ are $6^{-}$-vertices, then again $d_{2}(v)\leq 20$ by Lemma 2.2, and the same replacement as above yields a graph that is proper with respect to $G$. Thus $v$ must have a second $7(6)$-neighbour.
$(b)$ Suppose $m_{4}(v)=3$. We first show that $v$ has two $7$-neighbours. Assume three neighbours of $v$ are $6^{-}$-vertices. Then $d_{2}(v)\leq 20$ by Lemma 2.2. Setting $G^{\prime}=G-v+\{v_{2}v_{3},v_{1}v_{4}\}$ produces a graph that is proper with respect to $G$, contradicting Remark 2.1. Thus $v$ has two $7$-neighbours, and since it cannot have any $7(7)$-neighbour, these neighbours must be $7(6^{-})$-vertices. ∎
@@
    If $m_{4}(v)=0$, then $v$ has either a $7(5^{-})$-neighbour and a $6(4^{-})$-neighbour, or two $7(6^{-})$-neighbours.
@@
    If $m_{4}(v)=1$, then $v$ has either two $7(6^{-})$-neighbours and two $6(4^{-})$-neighbours, or three $7(6^{-})$-neighbours.
@@
    If $m_{4}(v)=2$, then $v$ has either two $7(5^{-})$-neighbours and two $7(6)$-neighbours, or three $7(5^{-})$-neighbours.
@@
Let $f_{i}=v_{i}vv_{i+1}$ and $f_{j}=v_{j}vv_{j+1}$ be the two $3$-faces incident to $v$, with $i<j$ (see Figure 3).
Whenever $d_{2}(v)\leq 20$, we reduce $G$ as follows: If $(i,j)=(1,2)$, we set $G^{\prime}=G-v+\{v_{2}v_{4}\}$. If $(i,j)=(1,3)$, we set $G^{\prime}=G-v+\{v_{2}v_{3},v_{1}v_{4}\}$. In each case, $G^{\prime}$ is proper with respect to $G$, contradicting Remark 2.1. Therefore, throughout the proof we may assume $d_{2}(v)\geq 21$.
$(a)$ Let $m_{4}(v)=0$. Suppose first that $v$ has a $7(7)$-neighbour. Then $v$ cannot have two $6^{-}$-neighbours; otherwise $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. Thus $v$ has two $7(6^{-})$-neighbours.
Now suppose $v$ has no $7(7)$-neighbour. Since $v$ cannot have all neighbours of degree at most $6$, it follows that $v$ has at least one $7$-neighbour; let it be $v_{r}$. Assume that $v$ has exactly one $7(6^{-})$-neighbour (otherwise the conclusion already holds). Then all other neighbours of $v$ are $6^{-}$-vertices. If $v_{r}$ were a $7(6)$-vertex, or if $v$ had any $6(5^{+})$-neighbour, then again $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. Hence $v_{r}$ must be a $7(5^{-})$-vertex, and $v$ must have a $6(4^{-})$-neighbour.
[…]
@@
Figure 3. Two possible 33-faces $f_{i}=v_{i}vv_{i+1}$ and $f_{j}=v_{j}vv_{j+1}$ incident to a $4(2)$-vertex.
$(b)$ Let $m_{4}(v)=1$. We first claim that $v$ has at least two $7$-neighbours. Indeed, if three neighbours of $v$ were $6^{-}$-vertices, then $d_{2}(v)\leq 20$ (since $m_{4}(v)=1$), a contradiction. Note also that $v$ has at most one $7(7)$-neighbour, because $m_{3}(v)=2$. We now distinguish cases depending on how many of the edges $v_{i}v_{i+1}$ and $v_{j}v_{j+1}$ lie in two $3$-faces.
Suppose that both $v_{i}v_{i+1}$ and $v_{j}v_{j+1}$ lie in two $3$-faces. Then $v$ cannot have any $6^{-}$-neighbour (otherwise $d_{2}(v)\leq 20$). Hence all its neighbours are $7$-vertices, and thus $v$ has three $7(6^{-})$-neighbours.
Suppose next that exactly one of $v_{i}v_{i+1}$, $v_{j}v_{j+1}$ lies in two $3$-faces. Then $v$ cannot have any $7(7)$- or $6(6)$-neighbour, so it has exactly two $7(6^{-})$-neighbours, say $v_{1}$ and $v_{2}$. Consider $v_{3}$ and $v_{4}$. If one is a $5^{-}$-vertex, or if both are $6$-vertices, then $d_{2}(v)\leq 20$, a contradiction. Thus one of $v_{3},v_{4}$ is a $7(6^{-})$-vertex and the other is a $6^{+}$-vertex. Hence $v$ has at least three $7(6^{-})$-neighbours.
Finally suppose that neither $v_{i}v_{i+1}$ nor $v_{j}v_{j+1}$ lies in two $3$-faces. Then $v$ has no $7(7)$-, $7(6)$-, $6(6)$-, or $6(5)$-neighbour. Thus $v$ has two $7(5^{-})$-neighbours, say $v_{1}$ and $v_{2}$. For $v_{3},v_{4}$: If one is a $5^{-}$-vertex, then the other must be a $7(5^{-})$-vertex (else $d_{2}(v)\leq 20$). If neither is a $5^{-}$-vertex, then either one is a $7(5^{-})$-vertex or both are $6(4^{-})$-vertices. This completes part (b).
$(c)$ Let $m_{4}(v)=2$. If $v$ had two $6^{-}$-neighbours, or a $7(7)$-neighbour, or a $6(6)$-neighbour, then $d_{2}(v)\leq 20$ (since $m_{4}(v)=2$), a contradiction. Thus $v$ has exactly three $7(6^{-})$-neighbours.
If $v$ had three $7(6)$-neighbours, then both $v_{i}v_{i+1}$ and $v_{j}v_{j+1}$ would have to lie in two $3$-faces, forcing again $d_{2}(v)\leq 20$, a contradiction. Thus $v$ has at most two $7(6)$-neighbours.
If $v$ has a $7(6)$-neighbour, then at least one of $v_{i}v_{i+1}$ or $v_{j}v_{j+1}$ lies in two $3$-faces, say $v_{i}v_{i+1}$. This forces that $v_{i+1}$ is a $7(6^{-})$-vertex, and all neighbours of $v$ other than $v_{i},v_{i+1}$ are $7(5^{-})$-vertices (otherwise $d_{2}(v)\leq 20$). Hence $v$ has either: two $7(6)$-neighbours and two $7(5^{-})$-neighbours, or three $7(5^{-})$-neighbours. ∎
@@
    If $m_{4}(v)=0$, then $v$ has one of the following: four $7(5^{-})$-neighbours, or three $7(5^{-})$- and one $7(6)$-neighbours, or three $7(5^{-})$- and one $6(4^{-})$-neighbours, or two $7(5^{-})$- and two $7(6)$-neighbours.
@@
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[3]$. If $d_{2}(v)\leq 20$, then $G^{\prime}=G-v+\{v_{1}v_{4}\}$ yields a graph that is proper with respect to $G$, contradicting Remark 2.1. Therefore, throughout the proof we may assume $d_{2}(v)\geq 21$. Note that at most one of the edges $v_{1}v_{2}$, $v_{2}v_{3}$, $v_{3}v_{4}$ is contained in two $3$-faces, and so $v$ has no $6(6)$- or $7(7)$-neighbour. Observe also that $v$ has no $5^{-}$- or $6(5)$-neighbour; $v$ has at most one $6$-neighbour; and $v$ has at most two $7(6)$-neighbours.
$(a)$ Let $m_{4}(v)=0$. As noted above, $v$ has at most one $6(4^{-})$-neighbour. If $v$ has such a $6(4^{-})$-neighbour, then it cannot have a $7(6)$-neighbour, since otherwise $d_{2}(v)\leq 20$, a contradiction. Thus all other neighbours of $v$ are $7(5^{-})$-vertices. If $v$ has no $6(4^{-})$-neighbour, then the only possible neighbourhood types are exactly those listed in the statement, since $v$ has at most two $7(6)$-neighbours.
$(b)$ Let $m_{4}(v)=1$. If $v$ has a $7(6^{+})$- or a $6$-neighbour, then $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. Hence every neighbour of $v$ must be a $7(5^{-})$-vertex. ∎
@@
    $v$ cannot have two $5(5)$-neighbours.
@@
    If $v$ has a $5(5)$-neighbour, then $v$ has no $4$-neighbour.
@@
    If $v$ has a $7(6)$-neighbour, then $v$ is adjacent to at most one $4$-vertex.
@@
    If $m_{4}(v)=0$, then $v$ has at least two neighbours that are neither $5^{-}$- nor $6(6)$-vertices.
@@
    If $m_{4}(v)=1$, then $v$ has at least three $6^{+}$-neighbours different from $6(6)$-vertex.
@@
    If $m_{4}(v)=1$ and $v$ has a $7(7)$-neighbour, then $v$ has at least four $6^{+}$-neighbours different from $6(6)$-vertex.